What’s the difference between Q (amount of heat energy) and delta Q? I’ve read some books according to which they’re the same thing.

Grade:

1 Answers

Utkarsh Shukla
9 years ago
See, Q means amount of heat energy, while ∂Q means the change in the amount of heat energy. For example,
if Q1=10J and Q2=13J, then
∂Q=Q2-Q1=13J-10J=3J